# XRZoo: A Large-Scale and Versatile Dataset of Extended Reality (XR) Applications

### v1

@@ I Introduction
The increasing importance of XR applications makes it urgent to conduct software engineering (SE) research, facilitating the reliability and efficiency of the whole software process from design, implementation, testing, and maintenance [6]. Many researchers have been drawn to address XR software challenges across diverse domains, such as automated interactive testing [7, 8], bug detection [9, 10], and static/dynamic analysis (e.g., for privacy auditing) [11, 12]. Similar to research in conventional software domains, the primary obstacle lies in subject collection. Due to the fragmentation issues of the dispersed XR ecosystems, diversified development technology stacks and competing hardware platforms, the platform-specific restrictions on accessing XR applications and associated data hinder the ability to collect comprehensive datasets, making the lack of large-scale, comprehensive, accessible, and high-quality datasets that span the diverse array of XR platforms and application categories. As a result, researchers often resort to reusing limited datasets (has risks of being outdated) or collecting small, non-representative samples. Such approaches can lead to biased analyses and experiments that are difficult to reproduce, thereby limiting the generalizability and impact of research in this domain.

### v2

@@ I Introduction
The increasing importance of XR applications makes it urgent to conduct software engineering (SE) and security research, facilitating the reliability and efficiency of the whole software process from design, implementation, testing, and maintenance [6, 7, 8, 9]. Many researchers have been drawn to address XR software challenges across diverse domains, such as automated interactive testing [10, 11], bug detection [12, 13], and static/dynamic analysis (e.g., for privacy auditing or clone detection) [14, 15].
Similar to research in conventional software domains, the primary obstacle lies in subject collection. To the best of our knowledge, there doesn’t exist a large-scale XR application dataset in the research community. The fragmented nature of dispersed XR ecosystems, coupled with diverse development technology stacks and competing hardware platforms, creates platform-specific restrictions on accessing XR applications and their associated data. This fragmentation severely limits the ability to collect comprehensive datasets, resulting in such lack of accessible, high-quality datasets that span the broad spectrum of XR platforms and application categories. As a result, researchers often resort to reusing limited datasets (has risks of being outdated) or collecting small, non-representative samples. Such approaches can lead to biased analyses and experiments that are difficult to reproduce, thereby limiting the generalizability and impact of research in this domain.